OpenStudy (anonymous):

How many inches thick is an arterial wall that measures 0.470um ?

sam (.sam.):

\[0.4700 \text{$\mu $m} \times\frac{1~m}{1\times 10^{6}\mu m}\times \frac{39.37~inch}{1~m}\] \[0.4700 \cancel{\mu m} \times\frac{1~\cancel{m}}{1\times 10^{6}\cancel{\mu m}}\times \frac{39.37~inch}{1~\cancel{m}}=1.85\times 10^{-5}\text{ }\text{inches}\]